# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v1

@@ Appendix A Supplementary proofs
Combining these upper bounds, we arrive at:
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]\leq\frac{16(2^{r_{2}}+1)(E_{1}+E_{2})H(2r)+9E_{3}H(s)1\{s<4\}}{(M_{0}2^{\ell})^{r_{2}}}.$
For the second required upper bound, we follow from (36) that
$\left\|\Delta\rho_{\phi,\ell}\right\|_{2}$ $\leq\|F^{(a)}_{\ell-1}\|_{2}|S^{(a)}_{\ell-1}|+\|F^{(b)}_{\ell-1}\|_{2}|S^{(b)}_{\ell-1}|+2\|F_{\ell}\|_{2}|S_{\ell}|$ (41)
$+\left\|\frac{\nabla Z_{\phi}(x,\phi)}{Z(x,\phi)}\right\|_{2}\left((S^{(a)}_{\ell-1})^{2}+(S^{b}_{\ell})^{2}+2(S_{\ell})^{2}\right),$ (42)
the same decomposition on event $A$ is also studied. We apply Lemma 3.1 and (39) to derive an upper bound for the second row. For the first row, it would be sufficient to find an upper bound for $\mathbb{E}[\|F_{\ell}\|_{2}|S_{\ell}|1_{\mathcal{A}^{c}}]$, applying Holder’s inequality to yield that
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}\left|S_{\ell}\right|1_{A^{c}}\right]\leq\left(\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\right]\right)^{1/2}\left(\mathbb{E}\left[\left|S_{\ell}\right|^{2}\right]\right)^{1/2}.$
We then utilize Lemma 3.1 to have
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}1_{\mathcal{A}^{c}}\right]\leq\frac{2d^{1/2}M_{\max}C_{2}}{M_{\ell}}H(2),\quad\mathbb{E}[\left|S_{\ell}\right|^{2}1_{\mathcal{A}^{c}}]\leq\frac{C_{2}}{M_{\ell}}H(2).$
Therefore the upper bound for $\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}\left|S_{\ell}\right|1_{A^{c}}\right]$ is $\sqrt{2}d^{1/4}M^{1/2}_{\max}C_{2}H(2)/M_{\ell}$. Finally we have the following upper bound for $\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|\right]$
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|\right]\leq\frac{(6\sqrt{2}d^{1/4}M^{1/2}_{\max}+4d^{1/2}M_{\max})C_{2}H(2)}{M_{\ell}}.$
A.2 Proof for Theorem 3.4
Theorem A.2.
Under the assumptions of Theorem 3.3, we have
$\mathrm{Var}\left[\rho_{\phi,M_{\ell}}\right]$ $\leq\frac{(16dM_{\max}^{2}+32d^{1/2}M_{\max})C_{2}H(2)}{M_{0}2^{\ell}}$
$+\frac{\left(8dM^{2}_{\max}+24\sqrt{2}d^{3/4}M^{3/2}_{\max}+36d^{1/2}M_{\max}\right)C^{2}_{2}H^{2}_{2}}{(M_{0}2^{\ell})^{2}}.$ (43)
Proof.
By Jensen’s equality, we have
$\mathrm{Var}\left[\rho_{\phi,M_{\ell}}\right]$ $=\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]\right\|_{2}^{2}\right]$
$=\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}+\rho_{\phi}-\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]\right\|_{2}^{2}\right]$
$\leq 2\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\right]+2\mathbb{E}\left[\left\|\rho_{\phi}-\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]\right\|_{2}^{2}\right].$ (44)
For the first term of (43), we begin with
$\rho_{\phi,M_{\ell}}-\rho_{\phi}=$ $\left(-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log\hat{Z}_{M_{\ell}}(x,\phi)\right)$
$-\left(-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log Z(x,\phi)\right)$
$=$ $\ \frac{\nabla g_{\phi,M_{\ell}}(x)}{g_{\phi,M_{\ell}}(x)}-\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)},$
and similarly, we consider the event
:=$\mathcal{B}:=\left\{\left|\frac{g_{\phi,M_{\ell}}(x)}{Z(x,\phi)}-1\right|>\frac{1}{2}\right\},$
and the similar decomposition used for event $\mathcal{A}$ in the proof of Theorem 3.3 is studied. Firstly we derive a rough bound for $\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}$ with (32)
$\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\leq 2\left\|\frac{\nabla g_{\phi,M_{\ell}}(x)}{g_{\phi,M_{\ell}}(x)}\right\|_{2}^{2}+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}^{2}\leq 4dM_{\max}^{2}.$
Then for event $\mathcal{B}$, by Lemma 3.1 again, we obtain
$\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\mathbf{1}_{B}\right]\leq 4dM_{\max}\mathbb{P}[B]\leq\frac{2^{s+2}dM_{\max}C_{s}}{(M_{0}2^{\ell})^{s/2}}H(s).$
On event $\mathcal{B}^{c}$, we have
$\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|^{2}_{2}$ $=\left\|\frac{\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}+\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\left(\frac{g_{\phi,M_{\ell}}(x)}{Z(x,\phi)}-1\right)\right\|^{2}_{2}$
$\leq 8\left\|F_{\ell}\right\|_{2}^{2}+8\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|^{2}_{2}S_{\ell}^{2},$
and
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}1_{\mathcal{B}^{c}}\right]\leq\frac{2d^{1/2}M_{\max}C_{2}}{M_{\ell}}H(2),\quad\mathbb{E}[\left|S_{\ell}\right|^{2}1_{\mathcal{B}^{c}}]\leq\frac{C_{2}}{M_{\ell}}H(2).$
These result in the following upper bound
$\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\right]\leq\frac{(16d^{1/2}M_{\max}+8dM_{\max}^{2})C_{2}H(2)}{M_{0}2^{\ell}}.$
For the second term of (43), we have
$\left\|\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]-\rho_{\phi}\right\|_{2}=\left\|\sum_{\ell^{\prime}=\ell+1}^{\infty}\mathbb{E}\left[\Delta\rho_{\phi,\ell^{\prime}}\right]\right\|_{2}\leq\sum_{\ell^{\prime}=\ell+1}^{\infty}\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell^{\prime}}\right\|_{2}\right],$
then applying Theorem 3.3 to yield the desired result
$\left\|\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]-\rho_{\phi}\right\|_{2}\leq\frac{(3\sqrt{2}d^{1/4}M^{1/2}_{\max}+2d^{1/2}M_{\max})C_{2}H(2)}{M_{0}2^{\ell}}.$ (45)
A.3 Proof of Theorem 3.6
Proof.
Applying Theorem 3.3, we have the following upper bound for the variance of the gradient estimators of GRR
$\mathrm{Var}\left[V_{\mathrm{GRR}}\right]$ $=\mathrm{Var}\left[\rho_{\phi,M_{{\underline{m}}}}\right]+\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\sum_{\ell^{\prime}={\underline{m}}+1}^{\ell}(1-p)^{-2\ell^{\prime}}\mathrm{Var}\left[\Delta\rho_{\phi,\ell^{\prime}}\right]$
$\leq\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\sum_{\ell^{\prime}={\underline{m}}+1}^{\ell}(1-p)^{-2\ell^{\prime}}B_{3}2^{-r_{2}\ell^{\prime}}$
$=\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+B_{3}\left(\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\frac{2^{(2\alpha-r_{2})({\underline{m}}+1)}(1-2^{(2\alpha-r_{2})(\ell-{\underline{m}})})}{1-2^{2\alpha-r_{2}}}\right)$
$=\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+B_{3}\left(\sum_{\ell={\underline{m}}+1}^{\infty}(1-p)^{\ell}p\frac{2^{(2\alpha-r_{2})({\underline{m}}+1)}-2^{(2\alpha-r_{2})(\ell+1)}}{1-2^{2\alpha-r_{2}}}\right)$
$=\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+B_{3}\left(\frac{1-2^{-\alpha}}{1-2^{2\alpha-r_{2}}}\left(\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{-\alpha}}-\frac{2^{2\alpha-r_{2}}2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right)\right)$
$\leq\frac{B_{1}+B_{2}}{2^{\underline{m}}}+B_{3}\left(\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right)$
$\leq\frac{B}{2^{\underline{m}}}+B\left(\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right),$
where $B=\max\{B_{1}+B_{2},B_{3}\}$ with $B_{1},\ B_{2}$ and $B_{3}$ representing the constants in the upper bounds in Theorems 3.4 and 3.3. The proof for the upper bound of $\mathrm{Var}[U_{\mathrm{GRR}}]$ follows a similar procedure. The average cost is
$=M_{0}2^{{\underline{m}}}+M_{0}(1-2^{-\alpha})\left(\frac{2\times 2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{(1-\alpha)}}-\frac{2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{-\alpha}}\right)$
$=M_{0}2^{\underline{m}}+M_{0}\frac{2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{(1-\alpha)}},$
completing the proof. ∎

### v2

@@ 6 Concluding remarks
Similarly, $\mathbb{E}[(S_{\ell-1}^{(a)})^{4}\mathbf{1}_{\mathcal{A}^{c}}]$ and $\mathbb{E}[(S_{\ell-1}^{(b)})^{4}\mathbf{1}_{\mathcal{A}^{c}}]$ have an upper of bound $\mathcal{O}(M_{\ell}^{-r_{2}})$. Combining these upper bounds, we arrive at
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]=\mathcal{O}(M_{\ell}^{-s/2})+\mathcal{O}(M_{\ell}^{-r_{2}})=\mathcal{O}(M_{\ell}^{-r_{2}}).$
The same decomposition on event $\mathcal{A}$ is also studied for $\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\right]$, yielding
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\right]=\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\mathbf{1}_{\mathcal{A}}\right]+\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\mathbf{1}_{\mathcal{A}^{c}}\right].$ (49)
Using (38) and (A) with $u=2$, we have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\mathbf{1}_{\mathcal{A}}\right]=\mathcal{O}(M_{\ell}^{-1}).$
It follows from (41) that on event $\mathcal{A}^{c}$,
$\left\|\Delta\rho_{\phi,\ell}\right\|_{2}$ $\leq\|F^{(a)}_{\ell-1}\|_{2}|S^{(a)}_{\ell-1}|+\|F^{(b)}_{\ell-1}\|_{2}|S^{(b)}_{\ell-1}|+2\|F_{\ell}\|_{2}|S_{\ell}|$
$+\left\|\frac{\nabla Z_{\phi}(x,\phi)}{Z(x,\phi)}\right\|_{2}\left((S^{(a)}_{\ell-1})^{2}+(S^{(b)}_{\ell-1})^{2}+2(S_{\ell})^{2}\right).$ (50)
For the upper bound of $\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\mathbf{1}_{\mathcal{A}^{c}}\right]$, we can similarly apply Lemma A.1 to derive an upper bound of $\mathcal{O}(M_{\ell}^{-1})$ for $\mathbb{E}\left[(S^{(a)}_{\ell-1})^{2}\mathbf{1}_{\mathcal{A}^{c}}\right]$, $\mathbb{E}\left[(S^{(b)}_{\ell-1})^{2}\mathbf{1}_{\mathcal{A}^{c}}\right]$ and $\mathbb{E}\left[(S_{\ell})^{2}\mathbf{1}_{\mathcal{A}^{c}}\right]$. Applying Hölder’s inequality and using (47) and (48) with $u=2$ yield
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}\left|S_{\ell}\right|\mathbf{1}_{\mathcal{A}^{c}}\right]\leq\left(\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\right]\right)^{1/2}\left(\mathbb{E}\left[S_{\ell}^{2}\right]\right)^{1/2}=\mathcal{O}(M_{\ell}^{-1}).$
Finally, we have $\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\right]=\mathcal{O}(M_{\ell}^{-1}).$
Appendix B Normalizing flows as a density estimator
Normalizing flow model maps noise $\mathbf{x}\sim\pi$ to data $\mathbf{y}$ through an invertible and differentiable transformation $f$, that is $\mathbf{y}=f(\mathbf{x})$. Then the probability density of $\mathbf{y}$ under the flow is obtained with a change of variables:
$p(\mathbf{y})=\pi(f^{-1}(\mathbf{y}))\left|\mathrm{\det}\left(\frac{\partial f^{-1}}{\partial\mathbf{y}}\right)\right|.$
Therefore, it is determined by the bijection $f$ and its inverse $f^{-1}$, which are often implemented by composing a series of invertible neural-network modules. Neural spline flows [13] used in our paper is built upon coupling layers, it maps $\mathbf{x}$ to $\mathbf{y}$ by using monotonic rational-quadratic splines, where each interval is defined by a monotonic rational-quadratic function taking the form of a ratio of two quadratic polynomials. These functions are easy to differentiate and, when restricted to monotonic segments, they are also analytically invertible. Moreover, both the derivative and the height at each knot are strictly more flexible. The spline itself maps the interval $[-B,B]$ to $[-B,B]$, and outside this range, the transformation is defined as the identity, resulting in a linear ‘tail’ that allows the overall transformation to accept unconstrained inputs.
The spline uses $K$ different rational-quadratic functions, with boundaries set by $K+1$ spline knots $\{u^{(k)},v^{(k)}\}_{k=0}^{K}$, with boundaries given by the coordinates $(u^{(0)},v^{(0)})=(-B,-B)$ and $(u^{(K)},v^{(K)})=(B,B)$. The remaining knots must satisfy the monotonicity conditions
$u^{(k)}<u^{(k+1)},\quad v^{(k)}<v^{(k+1)}.$
We specify the spline using $K-1$ arbitrary positive derivative values $\{\delta^{(k)}\}_{k=1}^{K-1}$ at the $K-1$ interior knots $(u^{(k)},v^{(k)})_{k=1}^{K-1}$, and set the boundary derivatives to $\delta^{(0)}=\delta^{(K)}=1$ in order to meet the linear tails.
The coupling transformation $\Psi$ maps the input $\mathbf{x}\in\mathbb{R}^{D}$ to the output $\mathbf{y}\in\mathbb{R}^{D}$ as follows:
  1. 1.
    Split the input $\mathbf{x}$ into two parts, $\mathbf{x}=[\mathbf{x}_{1:D^{\prime}-1},\mathbf{x}_{D^{\prime}:D}]$, where $D^{\prime}<D$.
  2. 2.
    A neural network is used to take $\mathbf{x}_{1:D^{\prime}-1}$ as input and output an unconstrained parameter vector $\phi_{i}$ of length $13K-$ 1 for each $i=D^{\prime},\ldots,D$.
  3. 3.
    Compute $\mathbf{y}_{i}=\Gamma_{\phi_{i}}(\mathbf{x}_{i})$ in parallel for $i=D^{\prime},\dots,D$, where $\Gamma_{\phi_{i}}$ is an invertible function parameterized by $\phi_{i}=[\phi_{i}^{w},\phi_{i}^{h},\phi_{i}^{d}]$, where $\phi_{i}^{w}$ and $\phi_{i}^{h}$ are of length $K$, and $\phi_{i}^{d}$ is of length $K-1$, vectors $\phi_{i}^{w}$ and $\phi_{i}^{h}$ are used to obtain the widths and heights of $K$ intervals. The cumulative sums of these widths and heights define the $K+1$ knot points $\{(u^{(k)},v^{(k)})\}_{k=0}^{K}$. Vector $\phi_{i}^{d}$ is obtained the derivative values $\{\delta^{(k)}\}_{k=1}^{K-1}$ at the interior knots.
  4. 4.
    Set $\mathbf{y}_{1:D^{\prime}-1}=\mathbf{x}_{1:D^{\prime}-1}$ and return $\mathbf{y}=[\mathbf{y}_{1:D^{\prime}-1},\mathbf{y}_{D^{\prime}:D}]$.
For the monotonic rational-quadratic function in the $k^{\mathrm{th}}$ bin, $w^{(k)}=u^{(k+1)}-u^{(k)}$ denotes the width of the interval, and $s^{(k)}=(v^{(k+1)}-v^{(k)})/w^{(k)}$ denotes the slope of the line connecting the coordinates. For $u\in[u^{(k)},u^{(k+1)}]$, the monotone invertible mapping is defined by
$\Gamma_{\phi_{i}}(u)=v^{(k)}+\frac{(v^{(k+1)}-v^{(k)})[s^{(k)}\xi^{2}+\delta^{(k)}\xi(1-\xi)]}{s^{(k)}+[\delta^{(k+1)}+\delta^{(k)}-2s^{(k)}]\xi(1-\xi)},$
where $\xi=(u-u^{(k)})/w^{(k)}\in[0,1]$, $k=0,\dots,K-1$.
A
B
C
Figure 4: Approximated posterior for nested APT and atomic APT. A. Two-moon model, from left to right: available ground truth, atomic APT with inner samples $M=100$, nested APT with inner samples $M=100$. B. Lotka-Volterra model, from left to right: ground truth simulated with SMC-ABC [4], atomic APT with $M=100$, nested APT with $M=100$. C. M/G/1 queue model, the setting are the same with Lotka-Volterra. We have plotted the scatter plot for each of the 2D sub-spaces. The histogram for each $\theta_{i},i=1,\dots,4$ is plotted on the diagonal with the ground truth parameter marked with dotted lines.
Appendix C Visualization of nested APT and atomic APT
Under the settings in Section 5, Figure 4 compares the approximated posteriors for nested APT and atomic APT. Both methods yield a comparable posterior estimation to the benchmark SMC-ABC. A noteworthy advantage of the proposed nested APT method is that it enables the application of a series of existing results for an optimizer with a biased gradient.
Appendix D Ablation studies for $\alpha$
Figure 5 shows ablation studies of $\alpha$ for TGRR-MLMC, where $\alpha=1.673$ is our proposed optimal value, $\alpha=1.8$ is the value that minimizes the expected cost. We observe that $\alpha=1.673$ exhibits a slightly better performance than $\alpha=1.8$ for the Lotka-Volterra model.
Figure 5: Ablation studies of $\alpha$ for TGRR-MLMC, where $\alpha=1.673$ is our proposed optimal value, $\alpha=1.8$ is the value that minimizes the expected cost.
Appendix E Inference on high-dimensional neuron models
We modeled a single-compartment Hodgkin-Huxley type neuron based on the equations in [49], where the dynamics of the membrane potential $V$ and channel gating variables $q\in\{m,h,n,p\}$ are governed by:
$C_{m}\frac{dV}{dt}$ $=g_{l}(E_{L}-V)+g_{Na}m^{3}h(E_{Na}-V)+g_{K}n^{4}(E_{K}-V)$
$\quad+g_{M}p(E_{K}-V)+I_{\text{inj}}+\sigma\eta(t),$
$\frac{dq}{dt}$ $=q_{\infty}(V)-q/\tau_{q}(V),\quad q\in\{m,h,n,p\},$
where $C_{m}$ is the membrane capacitance, $g_{L}$ is the leak conductance, $E_{L}$ is the leak reversal potential, and $g_{c}$ is the density of channels for ion types $c\in\{Na^{+},K^{+},M\}$. The gating variables $m,h,n,p$ govern the activation and inactivation of the respective channels, and their kinetics are dependent on the membrane potential $V$. The injected current $I_{\text{inj}}$ and Gaussian noise $\sigma\eta(t)$ drive the neuron dynamics. The steady-state functions $q_{\infty}(V)$ and the time constants $\tau_{q}(V)$ characterize the gating kinetics, with additional parameters controlling the spike threshold $V_{T}$ and the time constant of adaptation $\tau_{p}(V)$.
Following [23], we set $E_{Na}=53\,\text{mV}$ and $E_{K}=-107\,\text{mV}$. The parameters we aim to infer include: the maximal conductances $g_{Na},g_{K},g_{l},g_{M}$, the membrane threshold $V_{T}$, the leak reversal potential $E_{l}$, the maximum time constant $t_{\max}$, and the noise parameter noise, corresponding to the vector $\theta\in\mathbb{R}^{8}$ in the code implementation.
Figure 6: The comparison between atomic APT and TGRR-MLMC on Hodgkin-Huxley, measured with LMD and NLOG since ground truth posterior samples are inaccessible.
Figure 6 presents the LMD and NLOG values of the posterior distributions derived from the atomic APT and TGRR-MLMC methods over 10 rounds (10,000 simulations per round). In this high-dimensional task, we observe that the MLMC-based approach is able to improve performance, while atomic APT yields larger LMDs. However, we also find that MLMC requires more computational time due to the gradient calculation at multiple levels. The computational cost measured in minutes is $205.77\pm 18.38$ for atomic APT, and $379.10\pm 48.71$ for TGRR-MLMC.